The faces of a weighted cube are numbered 1,2,3,4,5, and 6, with one digit per face. When tossed, the cube is twice as likely to land with an even digit showing on its top face as an odd. Each odd number is as likely to appear as any other odd number, and each even number is as likely as any other even number. what is the probability that the cube, when tossed, shows a 5 on its top face? Express your answer as a common fraction.
Answer by Nate(3500)

even = e
odd = o
2o = e
e + o = 100% = 1
2o + o = 1
o = 1/3
Chances of getting an odd = 1/3
Chances of getting an even = 2/3
5 is odd ~> (1/3)
1,3,5 odd numbers ... 3 odds ~> 1/3 of getting a 5
(1/3)(1/3) = 1/9 or about 11.11%